OpenStudy (anonymous):

find the surface area obtained by rotating the curve y=e^x on [0,1] around the x-axis

OpenStudy (anonymous):

the surface area formula is AREA = integral 2pie*f(x)*sqrt(1+[f '(x)]^2)dx

OpenStudy (dumbcow):

f'(x) = e^x radius is e^x +1 \[\rightarrow 2\pi \int\limits_{0}^{1}(e^{x}+1)\sqrt{1+e^{2x}} dx\]

OpenStudy (anonymous):

how do you integrate that?

OpenStudy (dumbcow):

sorry about that, also i was wrong about the radius, its just e^x \[2\pi \int\limits_{0}^{1}e^{x}\sqrt{1+e^{2x}} dx\] substitute u = e^x du = e^x dx \[2\pi \int\limits\limits_{0}^{1}\sqrt{1+u^{2}} du\] substitute u = tan@ du = sec^2@ d@ \[2\pi \int\limits\limits_{0}^{1}\sec^{3} \theta d \theta\] use integration by parts u = sec dv = sec^2 du = sectan v = tan \[\rightarrow \int\limits_{}^{}\sec^{3}\theta = \sec \theta \tan \theta-\int\limits_{}^{}\sec \theta \tan^{2}\theta\] using identity tan^2 = 1-sec^2 \[\rightarrow \int\limits_{}^{}\sec^{3}\theta = \sec \theta \tan \theta+\int\limits_{}^{}\sec \theta-\int\limits_{}^{}\sec^{3} \theta \] \[\rightarrow \int\limits_{}^{}\sec^{3}\theta =\frac{1}{2} \sec \theta \tan \theta +\frac{1}{2}\int\limits_{}^{}\sec \theta \] \[\rightarrow \int\limits\limits_{}^{}\sec^{3}\theta =\frac{1}{2} \sec \theta \tan \theta +\frac{1}{2}\ln (\sec \theta+\tan \theta)\] Now notice tan@ = u = e^x sec@ = sqrt(1+tan^2) = sqrt(1+e^2x) Substitute back and evaluate \[2\pi \int\limits_{0}^{1}e^{x}\sqrt{1+e^{2x}} = \pi[e^{x}\sqrt{1+e^{2x}}+\ln (e^{x}+\sqrt{1+e^{2x}})] from 0->1\] \[=\pi[e\sqrt{1+e^{2}}+\ln (e+\sqrt{1+e^{2}})] - \pi[\sqrt{2}+\ln (1+\sqrt{2})]\] \[\approx 22.943\]